### v1

# LLMs in Biomedicine: A study on clinical Named Entity Recognition

Abstract
Large Language Models (LLMs) demonstrate remarkable versatility in various NLP tasks but encounter distinct challenges in biomedicine due to medical language complexities and data scarcity. This paper investigates the application of LLMs in the medical domain by exploring strategies to enhance their performance for the Named-Entity Recognition (NER) task. Specifically, our study reveals the importance of meticulously designed prompts in biomedicine. Strategic selection of in-context examples yields a notable improvement, showcasing $\sim 15-20\%$ increase in F1 score across all benchmark datasets for few-shot clinical NER. Additionally, our findings suggest that integrating external resources through prompting strategies can bridge the gap between general-purpose LLM proficiency and the specialized demands of medical NER. Leveraging a medical knowledge base, our proposed method inspired by Retrieval-Augmented Generation (RAG) can boost the F1 score of LLMs for zero-shot clinical NER. We will release the code upon publication.
LLMs in Biomedicine: A study on clinical Named Entity Recognition
Masoud Monajatipoor UCLA monajati@ucla.edu                        Jiaxin Yang UCLA yangjiaxin0821@ucla.edu                        Joel Stremmel Optum joel_stremmel@optum.com
Melika Emami Optum melika.emami@optum.com                        Ehsan Mohaghegh Optum ehsan.mohaghegh@optum.com                        Mozhdeh Rouhsedaghat USC rouhseda@usc.edu                        Kai-Wei Chang UCLA kwchang@cs.ucla.edu
1 Introduction
LLMs such as OpenAI’s GPT-4 have demonstrated exceptional capabilities across diverse tasks and domains, exhibiting remarkable performance Espejel et al. (2023); Dai et al. (2023); Dong et al. (2022). These models could have a revolutionary impact on healthcare, however, their integration into medical research and practice lags behind Zhou et al. (2023); Vaishya et al. (2023); Nori et al. (2023a). Acknowledging this gap, it is crucial to examine the unique challenges presented by the medical field that contribute to this discrepancy. Specifically, LLMs encounter challenges in medical Information Extraction (IE) Gutierrez et al. (2022); Moradi et al. (2021) due to the scarcity of high-quality clinical data in their pretraining, compounded by the requirement for a nuanced comprehension of the text for this task Gu et al. (2023). Medical entities can have multiple synonyms and abbreviations which makes them difficult for models to recognize Grossman Liu et al. (2021). Besides, context sensitivity takes on even greater significance in biomedicine compared to the general domain. The specificity of entity types and the complexity of their interrelations necessitate a level of background knowledge that standard prompts may fail to provide. LLMs are primarily exposed to vast amounts of generic text data limiting their effectiveness in managing the intricate nuances of medical language Kumari et al. (2023); Karabacak and Margetis (2023).
In this paper, we concentrate on NER, a foundational task for various applications in biomedicine. In our initial analysis, we broaden the scope of TANL Paolini et al. (2021) and DICE Ma et al. (2022), two text-to-text formats initially proposed for model training, adapting their use to prompt design specifically for the biomedical NER. Our findings reveal that the relative effectiveness of the resulting prompt pattern varies based on specific dataset characteristics. Subsequently, we investigate the importance of example selection in In-Context Learning (ICL) and demonstrate the value of nearest neighbor example selection using example encoders pretrained on biomedical text for the biomedical NER. A key question that arises in the deployment of LLMs concerns the comparative advantage of closed-source LLMs versus fine-tuned open-source ones. In our third study, we shed light on this question by presenting an assessment of performance and cost across various experiments. Furthermore, we explore the integration of external clinical knowledge to further refine LLM capabilities Gao et al. (2023c); Zakka et al. (2024). Leveraging the insights gained from these techniques, we present a novel data augmentation method incorporating a medical knowledge base, e.g., UMLS Bodenreider (2004), which substantially improves zero-shot clinical NER.
2 Background and Preliminaries
Prompt engineering
Prompt tuning White et al. (2023); Lester et al. (2021); Ding et al. (2021) as its own research field shows that skillfully crafted prompts can significantly enhance LLM understanding for complex tasks Lu et al. (2021); Kaddour et al. (2023); Webson and Pavlick (2021). Researchers have explored different formats to examine how LLMs handle IE tasks Wang et al. (2023c); Gutierrez et al. (2022); Wang et al. (2023b). Another type of prompting is ICL Brown et al. (2020), where LLMs use a limited set of "input-output" pairs within the prompt along with a query input as demonstrations of what the task output should be. In this realm, Liu et al. (2021); Min et al. (2022); Gao et al. (2023a) demonstrated that choosing targeted in-context examples over random sampling leads to more accurate model responses.
Named Entity Recognition
GPT-NER Wang et al. (2023b) was one of the first methods to incorporate a unique symbol to transform the sequence tagging task into text generation via ICL with GPT-3 Brown et al. (2020), achieving performance levels on par with fully supervised baselines. Following this work, Gutierrez et al. (2022); Moradi et al. (2021) have shown that LLMs are not skilled few-shot learners in the biomedical domain. However, recent advancements, such as GPT-4, have increased LLM performance on various tasks Tian et al. (2024); Hu et al. (2024a); Nori et al. (2023a) including in the biomedical domain Hu et al. (2024b). In the direction of knowledge distillation from LLMs Wang et al. (2023c); Gu et al. (2023), UniNER Zhou et al. (2023) presented a targeted distillation technique coupled with instruction tuning to develop an efficient open-domain NER model. Our research draws from these works and explores the capabilities of LLMs for clinical NER, employing prompt design, strategic ICL example selection, and data augmentation via an external knowledge base to enhance performance.
Problem definition
Assume the data sample is represented as $(X,Y)$ and the goal is to develop a model, denoted as $f:(X\times T)\rightarrow Y$, where X signifies the input set, T represents a predetermined collection of entity types, and Y denotes the set of entities types within the provided input. The task is to predict the entity type of each input word among the set T. We followed the standard practice of using the F1 score for evaluation purposes.
Datasets
We use three biomedical NER datasets with different entity types: I2B2 Uzuner et al. (2011) which includes test, treatment, and problem entities, NCBI-disease Doğan et al. (2014) consisting of the disease entity, and BC2GM Smith et al. (2008) containing the gene entity.
3 Influence of Input-Output Format
Recent studies demonstrated the importance of prompt engineering for various tasks Wang et al. (2023a); Gao et al. (2023b); Nori et al. (2023b). We studied the influence of input-output format by adapting TANL Paolini et al. (2021) and DICE Ma et al. (2022) formats for the biomedical NER. In TANL, the task is framed as a translation task which involves augmenting the text by tagging entity types for each word directly within the text. The method is exemplified in Fig 7, showcasing how the text incorporates entity types.
[…]
Figure 1: TANL input/output format for NER task.
Then, the generated output is decoded into the BIO format for the assessment. In the refined DICE format, the input-output format involves adding a description for each entity type in a template following DEGREE Hsu et al. (2021). Given an input text and corresponding labels, the desired output should be the input followed by the phrase "entity type is <entity_type>. <entity_description>. entity is <entity>" for each class label, e.g., test, treatment, and problem in the I2B2 dataset. Then, we expect the model to output the same template filling out the <entity> with the corresponding entities in the given text as demonstrated in Fig 2. For the entity types with no matched entities in the sentence, the output returns <entity> token in the output. Examples for the NCBI-disease and BC2GM datasets are presented in Appendix A.
[…]
Figure 2: DICE input/output format for NER task.
Our experiments in Table 1 reveal that neither format consistently outperforms the other, rather the effectiveness of each format varies depending on the complexity of the dataset and model size. To maintain consistency in the rest of our experiments, we opted for the TANL format, in which the input-output relationship exhibits a more straightforward pattern.
input-output format I2B2 NCBI-disease BC2GM
GPT-3.5-turbo DICE 50.0 62.0 55.6
TANL 59.7 51.3 50.8
GPT-4 DICE 70.1 77.8 67.9
TANL 73.5 70.0 69.6
Table 1: Comparison of TANL and DICE using GPT-3.5-turbo/GPT-4 on three benchmark datasets. The superiority of any single format varies with the complexity of the dataset and model size.
4 In-Context Examples Selection: A Key to Improving ICL Outcomes
In-context examples can be randomly chosen from the training set, however, researchers have demonstrated that the performance of ICL depends on the order and similarity of ICL examples to the test samples Liu et al. (2021); Min et al. (2022); Gao et al. (2023a); Zhang et al. (2023). Liu et al. (2021) presented Knn-Augmented in-conText Example selection (KATE). KATE identifies in-context examples selectively using nearest neighbor search on example embeddings, leading to better performance than random example selection. We tested KATE on TANL formatted examples with 16-shot ICL using four different LM encoders (w/o fine-tuning) to produce example embeddings. We used MPNET Song et al. (2020) for its popularity and performance on sentence embedding benchmarks111 https://www.sbert.net/docs/pretrained_models.html, SimCSE Gao et al. (2021) for its documented performance as an alternative to standard sentence transformers, and BioClinicalBERT Alsentzer et al. (2019) and BioClinicalRoBERTa Gururangan et al. (2020) for their dominance on clinical data tasks Lehman et al. (2023).
Our results summarized in Table 2 show that strategic in-context example selection via KATE outperforms random selection. BioClinicalRoBERTa achieved the best results among all example encoders tested. The strong performance of BioClinicalBERT and BioClinicalRoBERTa underscores the importance of using LM encoders pretrained on the biomedical text when applying KATE for the biomedical NER.
KATE method vs random sampling I2B2 NCBI-disease BC2GM
GPT-3.5-turbo Random sampling 59.7 51.3 50.8
[…]
BioClinicalBERT 78.9 78.8 72.0
[…]
SimCSE 76.1 76.1 68.8
Hu et al. (2024b) 49.3 - -
GPT4 Random sampling 73.5 70.0 69.6
[…]
BioClinicalBERT 88.1 88.0 79.4
[…]
SimCSE 86.6 86.5 77.9
Hu et al. (2024b) 59.3 - -
Table 2: Random example selection versus KATE with medical/non-medical encoders. KATE significantly outperforms random sampling in all settings, and LMs pretrained on the biomedical text outperform strong, general domain encoders.
[…]
Figure 3: An overview of Dictionary-Infused RAG
5 In-Context Learning or Fine-Tuning?
Within the scope of LLMs for biomedical applications, an essential question is whether to prompt a closed-source LLM via ICL or fine-tune an open-source one. On the other hand, comparing two different LLMs employing divergent strategies is not straightforward. To provide some insight into this dilemma, we examined two key factors, performance and cost, for clinical NER, and present a detailed analysis under various experiment settings. This comparison offers valuable perspective into the right strategy given the task and dataset attributes. For fine-tuning, we used LoRA Hu et al. (2021), and its details can be found in Appendix A. The cost of fine-tuning comes from training an LLM on a large labeled dataset while the cost of ICL mainly comes from calling the API for each input query. For 16-shot ICL experiments, we calculated the cost based on the number of processed and generated tokens considering the average text size based on the current LLM API pricing.222 https://openai.com/pricing The estimated cost for the entire test set of each benchmark dataset considering the input text, prompt, and generated text size using the TANL format is summarized in Table 3. Referring to the OpenAI API for fine-tuning pricing, we also estimated the cost for fine-tuning LLama2-7B which is summarized in Table 3. Interestingly, for the I2B2 dataset, GPT-3.5-turbo with a much cheaper cost outperforms fine-tuning Llama2-7B.
Language model I2B2 NCBI-disease BC2GM
Performance GPT-3.5-turbo w/ KATE 78.9 78.8 72.5
GPT4 w/ KATE 88.4 88.3 80.7
Llama2-7B 76.2 91.3 75.1
Cost GPT3.5-turbo w/ KATE \$0.35 \$0.11 \$1.34
GPT4 w/ KATE \$10.42 \$3.12 \$40.13
Llama2-7B \$5.23 \$5.24 \$5.69
Table 3: Analysis of ICL vs fine-tuning LLMs: assessing performance and cost implications. Fine-tuning LLama2 exhibits superior outcomes on NCBI-disease, whereas GPT-4, enhanced by KATE using a biomedical encoder, achieves more favorable results on both the I2B2 and BC2GM datasets.
6 Dictionary-Infused RAG
Retrieval-Augmented Generation (RAG) Lewis et al. (2020) is a technique to enhance the capabilities of LLMs by integrating external information or knowledge into the generation process. This method involves retrieving relevant documents from a large corpus and providing this external knowledge in the input context to improve the quality and relevance of the generated text. Inspired by RAG, we developed a new method, DiRAG, to utilize UMLS as an external resource to augment the input data for the clinical NER task. The process with detailed prompts is visualized in Fig 3. Unlike traditional RAG techniques that rely on embedding similarities to retrieve relevant documents, our approach initially employs GPT-3.5-turbo to tackle a more straightforward task: identifying all words within an input text that could potentially qualify as named entities. Then, we look up each selected word in an external knowledge base, e.g., UMLS to augment the input data with useful information such as definition. Finally, we call GPT-4 with augmented input text. A similar prompting format to UniNER is used to do a fair comparison. We tested this approach on zero-shot NER and compared it with baselines in Table 4. Our proposed strategy with GPT-4 outperforms the baseline on I2B2 and NCBI-disease datasets with an F1 score improvement of 5.6% and 0.5% respectively. Additionally, the approach enhanced the performance of GPT-3.5-turbo on both the I2B2 and NCBI-disease datasets by 10.4% and 6.9% respectively. It proved ineffective for the BC2GM dataset due to the nature of the UMLS knowledge base which is predominantly tailored to medical terminology rather than biogenetics. We expect our approach to outperform GPT-4 on BC2GM if a more relevant knowledge base is employed.
I2B2 NCBI-disease BC2GM
UniversalNER Zhou et al. (2023) 40.4 61.4 47.2
Rohanian et al. (2023) w/ GPT-3.5 - 33.4 32.0
Hu et al. (2024b) w/ GPT-3.5-turbo 43.6 - -
Hu et al. (2024b) w/ GPT-4 52.6 - -
[…]
GPT-3.5-turbo w/ DiRAG 43.55 45.0 28.8
[…]
GPT-4 w/ DiRAG 60.04 62.17 53.83
Table 4: Zero-shot clinical NER using GPT-3.5-turbo and GPT-4 w/ and w/o DiRAG in comparison to SOTA. Our method improved zero-shot NER significantly for I2B2 and NCBI-disease datasets.
7 Conclusion
In this paper, we explored LLMs by customizing various prompting techniques specifically for the biomedical NER. Through a detailed comparative analysis, we have highlighted the vital role of ICL and the selection of contextually pertinent examples, demonstrating the benefits to clinical NER tasks. Moreover, our investigation into incorporating external clinical knowledge resulted in a novel data augmentation approach, considerably advancing the capabilities of zero-shot clinical NER with LLMs.
Limitations
While we have shown the potential of enhancing LLM performance for clinical NER, the experiments in this paper are limited in two aspects mainly due to computational constraints. (1) TANL uses a straightforward text-to-text format while DICE uses additional descriptions. Future work could attempt to simplify DICE or combine it with TANL. Ablation studies on components of each format could help researchers design new prompt formatting strategies. (2) Our RAG-based method exclusively utilizes UMLS as the knowledge base which is limited in its vocabulary. For medical terms not covered by UMLS, we did not augment the input text, though other knowledge bases such as Wikipedia could potentially serve as an alternative.
8 Aknowledgment
This work was supported in part by Optum. We would like to thank the reviewers for their feedback to improve this research work.
Appendix A Appendix
Fine-tuning setting for Llama
We fine-tuned Llama2-7B for three epochs and maintained a batch size of 16, learning rate of 2e-4, and cap the maximum sequence length at 512, truncating any sequences that exceeded this limit. The LoRA dropout rate is adjusted to 0.1, and the LoRA $\alpha$ and rank parameters are also set at 16 and 32 respectively. The training was done on 4 NVIDIA Tesla V100 GPUs for an approximate of 2 days.
TANL and DICE more examples
In this section, we visualize two examples of each format for NCBI-disease and BC2GM datasets for demonstration.
[…]
Figure 4: TANL input-output format example for NCBI-disease dataset
[…]
Figure 5: DICE input-output format example for NCBI-disease dataset
[…]
Figure 6: TANL input-output format example for BC2GM dataset
[…]
Figure 7: DICE input-output format example for BC2GM dataset